The sum of three numbers is 18. The sum of the first and second numbers is 15. The first number is 3 times larger than the third number. Solve for the three variables
Answer by Maths68(1474)

Let
Third Number = x
Second Number = y
First Number = 3x
The sum of 3 numbers is 18
x+3x+y=18
4x+y=18............(1)
The sum of the first two numbers is 15
3x+y=15..............(2)
Take (2)
3x+y=15
y=15-3x.........(3)
Put the value of y in (1)
4x+y=18
4x+(15-3x)=18
4x+15-3x=18
x=18-15
x=3
Put the value of x in (2)
3x+y=15
3(3)+y=15
9+y=15
y=15-9
y=6
Third Number = x = 3
Second Number = y = 6
First Number = 3x = 3(3) = 9